If f(x)=X^2+2 and g(x)=2*x-1,
find f(g(1)).
Thanks,
Carolyn

Answer by venugopalramana(3286)

If f(x)=X^2+2
and g(x)=2*x-1,
g(1)=2*1-1=2-1=1
find f(g(1)). =f(1)=1^2+2=3


Thanks,
Carolyn